Microeconomics – Week #8 Assignment Trade Restrictions - Tariffs Directions: Answer the questions below based on the graph. 1. How many units will the domestic firms produce without trade? 2. How many units will the domestic firms produce without a tariff if the foreign producer can sell the product at a $4 price? 3. How many units will the foreign firms produce / sell if a government tariff of $2.00 is imposed on foreign goods? 4. What will be the total government revenues if a tariff of $2.00 is imposed on foreign goods? 5. What will be the total deadweight losses if a tariff of $2.00 is imposed on foreign goods? Price $11.00 $7.50 $6.00 $4.007 4 69 12 15 Quantity
Answer #1

1. The domestic eqm will be at 9 units. This will be at the intersection of demand and supply curve.

2. When the price is $4 the domestic firm will produce 4 units of the good. This will be given as the intersection of price and the supply curve. The foreign firm will supply 15-4 units = 11 units.

3. When the government imposes a tariff of $2, the price increases to $6, where the domestic firm is supplying 6 units and the demand is 12 units. So, the foreign firm will supply the remaining 12-6 units = 6 units.

4. Revenue= Tariff * quantity imported = $2 * 6 = $12

5. Deadweight loss will be the area of triangle. It is the loss in total surplus because of introduction of tariff as compared to no tariff case. DWL = 1/2 * 6 * ($7.5-$6) = 0.5 * 6 * 0.5 = 1.5 units